What is the specific heat capacity of a 24.93g metal that absorbs 476.87J of heat if it undergoes a 72.8°C temperature change?
1
Expert's answer
2019-11-22T04:34:41-0500

C = Q / (m * (t2-t1)) = 476.87J / (0.02493kg * 72.8°C) = 262.75 J/(kg °C).